# SARATR-X: A Foundation Model for Synthetic Aperture Radar Images Target Recognition

### v1

@@ 2 Related Work
As shown in Table I and II, visual foundation models are booming in remote sensing. Many models are proposed for various modalities and tasks. Our studies focus on the foundation model for SAR ATR, i.e., SAR images-based target classification and object detection. In the following, we introduce the recent development of remote sensing foundation models.
2.1 Foundation models in remote sensing
@@
Remote sensing foundation models [30] have received widespread attention in the last three years. Researchers have made many breakthroughs, achieving effective learning on various modalities and tasks. In terms of pre-training datasets, researchers use existing large-scale datasets or collect a large number of samples from different sources. As for the model backbone, researchers have improved attention mechanisms, positional encoding, and other aspects to enhance the perception of complex spatial information. MIM has been used to learn spatial-temporal contextual information, while contrast learning is applied to multi-modal learning.
@@
SatMAE [13] proposed a novel masking strategy and temporal and spectral positional encoding for multi-spectral and temporal images. With a new dataset fMoW Sentinel with 13 frequency bands, SatMAE achieved new performance on scene classification and semantic segmentation tasks. RVSA [77] improved the pretraining ViT backbone with a rotated varied-size window attention method for the arbitrary-oriented objects. This work shows the importance of learning targets’ complex spatial contextual relationships in remote sensing images. RingMo [65] used a patch incomplete mask strategy for dense and small objects. Based on their self-built 2 million images, RingMo has proven effective in many tasks. RingMo-Sense [90] offered a three-branch network and masking strategy to model the spatio-temporal interaction for temporal images. CMID [55] combined contrastive learning and masked image modeling to learn global semantic information and local spatial information. GFM [52] focused on the differences between natural and remote sensing images and employed a multi-objective continual pretraining approach to leverage both knowledge. DiffusionSat [32] was the first remote sensing generative that employs geographic information embedding in stable diffusion. Scale-MAE [58] reconstructed images at different frequencies with improved positional encodings of ViT. FG-MAE [79] employed different hand-designed features to replace the original pixels in the MIM and improve the feature quality. And SMLFR [17] used a low pass filter to eliminate high-frequency information from the image pixel.
@@
Multi-modal remote sensing foundation models have been developed, such as SkySense [26] and OFA-Net [87]. SkySense [26] proposed a multi-granularity contrastive learning method to learn representations of different modalities. Besides, a GEO-context prototype was applied to embed geographical contextual information. OFA-Net [87] applied a shared Transformer backbone for multiple modalities. Besides this, there are vision-language models [41], such as EarthGPT [98], SkyEyeGPT [93] and LHRS-Bot [54], which incorporate large language models to different remote sensing images modalities. However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only has 10,554 SAR ship images, much less than 84,838 infrared images and 907,945 optical images.
Therefore, there is relatively less research on SAR foundation models due to the scarcity and fragmentation of the high-resolution SAR target datasets in Fig 2. We want to explore the visual foundation model based on SAR images with target recognition to stimulate research enthusiasm in this direction.
Figure 2: Timeline of measured and simulated datasets for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization). SAR target datasets have increased rapidly since the 2020s, greatly enriching the diversity of targets, scenes, and sensors. Target categories are mainly fine-grained vehicles, ships, and aircraft. Mixed means that the target category includes more than one category, such as a mixture containing various vehicles, ships, and aircraft in IRIS-SAR. Scenes contain cities, harbors, airports, etc. Sensors are devices with different resolutions and bands under satellite or airborne platforms. However, the number of samples (No. samples) in most datasets is only a few thousand with several categories due to the high costs and difficult annotation. It inspired us to propose SARATR-X with an integrated dataset in Table III.
@@
TABLE III: Description of our pre-training dataset for SARATR-X, which contains 14 open-source SAR datasets for pre-training. # Img.: Number of images. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Pol.: Polarization. Large SAR imagery in detection datasets contains more target and scene types than the annotation. We cropped some large images to increase the training samples.
@@
Dataset Year # Img. Img. size # Target # Scene Res. (m) Band Pol. Description
@@
2.2 Foundation models in SAR
Researchers have explored SAR foundation models in different aspects, as shown in Table II. Inspired by previous studies, we systematically investigated how to construct a SAR ATR foundation model.
Early SSL was often used as a regularization loss for classification tasks. RotANet [82] predicted the rotational pattern of MSTAR vehicle targets to capture azimuthal features for the classification task. UACL [88] combined data augmentation and adversarial samples in contrastive learning to improve the model’s robustness to various adversarial attacks. PGIL [29] used contrastive learning between the SAR complex images’ sub-frequency feature and the amplitude images’ deep feature to inject physical knowledge into the classification task. Recently, SSL has been used in model pre-training and fine-tuning frameworks. BIDFC [92] proposed weakly contrastive learning for pre-training in fine-grained vehicle datasets MSTAR and used Gaussian noise data augmentation to simulate SAR image noise. TSCL [56] applied the pre-processing of SAR images before data augmentation in contrastive learning. FG-MAE [79] discussed different hand-craft features for multi-spectral and SAR images and used the HOG feature for SAR. Our previous study, SAR-JEPA [38] and MSFA [43], focused on target classification and object detection. SAR-JEPA [38] applied Local reconstruction and multi-scale gradient feature to capture target spital signatures better. MSFA [43] proposed a multi-stage with filter augmentation pre-training framework to use large-scale RBG and SAR data for detection.
Our insight. These studies have demonstrated that SSL can achieve performance improvements on various categories [38] and tasks [38, 43], and can even be comparable to the performance of specially designed supervised methods [92, 43]. These inspired us to conduct systematic research for foundation models to achieve general SAR target recognition, especially in this era of big data. Firstly, we must extend the pre-dataset with different tasks and scenarios based on existing research. Secondly, a suitable model backbone needs to be discussed for the small target characteristics of remote-sensing images. Thirdly, SSL needs high-quality guide signals from SAR images under noise interference. Finally, we need to evaluate the performance of foundation models comprehensively.
3 Approach
We aim to construct a foundation model for general ATR from large-scale SAR images via the SSL method. As described above, the increasing SAR datasets and SSL studies inspired us to build a foundation model for SAR ATR. Our approach revolves around pre-training datasets, model backbones, SSL methods, and evaluation tasks to provide a systematic benchmark for foundation models in SAR ATR.
Figure 3: Two steps of our pre-training. The first is to perform MIM on ImageNet to obtain better initialization weights, as shown in Fig 5 (c). Besides, we can easily obtain the open-source pre-training weights in step 1, thus reducing the pre-training time. The second is to perform MIM on SAR images with high-quality guide signals. We use multi-scale gradient features to suppress speckle noise and extract target edges.
3.1 Creating a Diverse Pre-training Dataset
@@
Existing work mainly used MSTAR [1] as a pre-training dataset. While MSTAR is a high-quality vehicle target slice dataset, its commonly used samples are only a few thousand. Besides, this dataset suffers from background bias due to a single imaging scene [39]. By comparison, the pre-training set ImageNet-1K in computer vision contains 1.4 million images with different categories and scenes. Using an insufficient SAR target pre-training dataset would underestimate the potential of this research area. Since diverse targets, scenes, and sensor conditions constitute a huge data sampling space in real-world situations, constructing a large pre-training dataset for the foundation model is central to unifying various SAR ATR tasks.
@@
The increasing SAR target datasets in Fig. 2 is a primary motivation for achieving this goal. Although SAR images are expensive and no single dataset contains all popular target categories and various imaging conditions, collecting target samples from various open-source datasets can still construct a pre-training dataset with different categories, scenes, and sensors. Therefore, we build a new pre-training dataset containing 186,600 SAR target samples from 14 open-source SAR target datasets in Table III. We construct the dataset to include, as much as possible, target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of different resolutions and bands).
3.2 Choosing a Scalable Model
@@
We consider two model backbones for SAR target recognition. For the first, we apply Vision Transformer (ViT) [18], commonly used in SSL and has good scalability in model parameters. For the second architecture, we experiment with ConvNeXt-V2 [83], which has the same scalability as ViT and maintains the efficiency of convolutional neural networks. In addition to the scalability of the model parameters, the image properties need to be considered for remote sensing. The SAR target usually has a small foreground and a dynamic context range. Our previous study, MSFA [43], finds that the Swin Transformer is better than ViT with a hierarchical structure, but it is unsuitable for drop patches in MIM to save computing resources. Therefore, we also consider a variant of ViT, Hierarchical Vision Transformer (HiVit) [99], which improves the input spatial resolution and retains the property of ViT with MIM.
3.3 Pre-Training with two steps
We use MIM as the pretext task and masked autoencoders (MAE) [27] to save computational resources with dropping patches. MIM can help the foundation model achieve SAR image interpretation by recognizing the contextual relationship around objects. Here is a key point when applying MIM. SAR images belong to coherent imaging, and values have speckle noise that can interfere with the pretext task. Therefore, SARATR-X uses two pre-training steps to build a foundation model in Fig 3.
@@ 3.3 Pre-Training with two steps
The first step is to perform MIM on ImageNet to obtain better initialization weights. We simplify the multi-stage pre-training of MSFA, which performs SSL on ImageNet with backbone, detection task pre-training on DOTA with the whole framework, and detection task fine-tuning on SAR images. SARATR-X uses the pre-training weight of ImageNet as initialization weights for the SAR pre-training step. This way enhances the diversity of attention during the SAR pre-training step in Fig. 5. In contrast, random initialization leads to a convergence of attention towards the same pattern in SAR pre-training with MAE. Besides, ImageNet pre-training backbone weights can be obtained from open source, thus greatly reducing pre-training time. We refer to using ImageNet pre-trained weights as SSL-ImageNet & SAR.
@@
The second step is to perform MIM with SAR images. As mentioned before, SAR image noise is a very tricky problem, and FG-MAR, SAR-JEPA, and MSFA have discussed many features, such as CannyEdge [8], HOG [14], CannyEdge [8], Haar-like [70], SAR-HOG [61] and SAR-SIFT [16]. We can use different feature combinations to get the best results, but here is the simplest gradient feature that follows our previous SAR-JEPA to avoid getting bogged down on complex feature selection. We use the Multi-scale Gradient Feature (MGF) [38] to suppress the speckle noise and extract the target shape.
@@
TABLE IV: The results on the classification dataset with 25 categories. We perform the linear probing setting on the few-shot SAR target classification. For the pre-training datasets, we recommend using ImageNet’s pre-trained weights as initialization before pre-training on SAR datasets. For the model backbones, we prefer the HiVit to recognize the small target in SAR images. Model backbones are their base version. The classification metric is average accuracy. Bold indicates the best result, underline is the next best result. SL. Supervised Learning. SSL. Self-Supervised Learning.
@@
setting dataset method 5 10 20
@@
Multi-scale gradient feature. The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in the SAR image. It means that speckle noise would cause the gradient computation to have false points in strong target regions. Previous studies [68, 6] have shown that the computing ratio is suitable for multiplicative noise. Here, MGF uses the gradient by ratio [61, 16] to obtain gradient features $G_{\text{m}}$.
$R_{i}$ $=\frac{M_{1}(i)}{M_{2}(i)},$ (1)
$G_{\text{H}}$ $=log(R_{1}),$ (2)
$G_{\text{V}}$ $=log(R_{3}),$ (3)
$G_{\text{m}}$ $=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}},$ (4)
where $R_{i}$ denotes the average ratio at different directions. $M_{1}(i)$ and $M_{2}(i)$ are the area averages on opposite sides of the current pixel along direction $i$. $i=1$ is the horizontal direction and $i=3$ means the vertical direction. The area averages can be computed from the input image and four fixed convolution kernels. Then, Eq. 2 and 3 use logarithms to solve the vertical gradient calculation [16]. $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient.
$\text{MGF}=\text{concat}(G_{\text{m1}},G_{\text{m2}},G_{\text{m3}})$ (5)
Due to the dynamic range required for various targets in remote sensing [42], MGF is constructed with convolutional kernels of different sizes. We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1.
3.4 Evaluating with Recognition Tasks
We merged fine-grained classification datasets of Vehicles, Ships, and Aircraft as a new SAR classification dataset named (SAR-VSA). This dataset is used for comparing the SSL model’s performance with a few-shot setting in Sec. 4. We then report SARATR-X results on existing classification and detection dataset settings with other methods in Sec. 5.
4 Experiments of SARATR-X
First, we perform the SSL on the pre-training dataset without label information. Then, we fine-tune the pre-trained model on SAR-VSA with few-shot classification task and linear probing setting to analyze the improvement of SARATR-X. Finally, we discuss the scalability of the proposed approach.
We perform pre-training on 8 NVIDIA RTX3090 GPUs. The SAR pre-training dataset consists of fourteen SAR datasets. The few-shot SAR classification dataset, SAR-VSA, contains 25 fine-grained targets from the three SAR datasets. It is difficult to ensure training convergence by fine-tuning the model’s whole parameters under small sample cases. Therefore, we use linear probing [27], which includes a batch normalization layer, to adjust the differences in the data statistical properties and reduce the number of fine-tuning parameters. Detailed settings are in the Appendix  A.
4.1 Comparison of Model Backbones
Table IV compares different model backbones for SAR ATR: ConvNeXt-V2 [83], ViT [18], and HiViT [99]. ConvNeXt-V2 and ViT represent the two main architectures: CNN and Transformer. HiViT combines Vit and Swin Transformer [47].
From the result, we can see that ViT outperforms ConvNeXt-V2. On the one hand, the ViT is more flexible than ConvNeXt-V2 in learning the contextual information in SAR images. On the other hand, multiple downsamplings in ConvNeXt-V2 result in the loss of small targets, while ViT maintains the same spatial resolution in different layers. HiVit, a visual Transformer with ViT’s flexibility and hierarchical representation, performs better than ViT and ConvNeXt-V2 in our self-supervised experiments. In particular, HiViT uses small (4 × 4) input patches, capturing small target features well. Fig. 5 demonstrates that HiViT has a better various attention distance than ViT due to the small target information in remote sensing.
4.2 Strategy of two-step pre-training
Here, we discuss the two-step pre-training strategy to make full use of the available model weights and SAR datasets. Table IV has four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR. SL-ImageNet is pre-training on ImageNet with supervised learning [^1] We used open-source weights from GitHub to conduct the experiments. The supervised weights of ConvNeXt-V2 and HiViT were obtained by supervised fine-tuning after SSL.; SSL-ImageNet is pre-training on ImageNet with SSL from scratch; SSL-SAR is pre-training on our SAR pre-training dataset from scratch; SSL-ImageNet & SAR pre-training the model on SAR dataset based on the initialized weight from SSL-ImageNet.
It is noticed that the additional supervised information introduced by SL-ImageNet does not necessarily improve the SAR ATR performance, e.g., the linear probing performance of SL-ImageNet for HiVit is lower than that of SSL-ImageNet. SSL-SAR can achieve better results than SSL-ImageNet using less data (12%), reflecting the huge differences in the target features between these two images. However, ImageNet pre-training weights can provide a good initialization for low features such as shape and texture in SSL with visible spectral remote sensing [67] and medical image [103]. Our experiments also confirmed this conclusion: using SSL-ImageNet as initialization weights improves the pre-training performance of SAR images in Table IV and attention diversity in Fig. 5. Therefore, our SARATR-X uses the SSL-ImageNet & SAR setting to complement the richness of the pre-training.
4.3 Design of target signal for SAR images
@@
TABLE V: Comparisons of different target features for MIM. We found many target features are unsuitable for the multiplicative speckle noise in SAR images. This result inspired us to discuss different gradient features for SAR SSL. The pre-training setting SSL-SAR and the model backbones are the base version. The classification metric is average accuracy. Bold indicates the best result, underline is the next best result.
@@
After discussing the backbone and strategy, we focus on the target features for SSL methods with SAR images. Due to the unique multiplicative speckle noise in SAR images, a key point for MIM is designing high-quality guide signals. As shown in Table V, we consider five target features (pixel value [27], low pass filter [46], HOG feature [79], deep feature [2], SAR-HOG [61], and gradient by ratio [16, 61]). All SSL methods use the SSL-SAR setting and base version.

### v2

@@ II Related Work
Visual foundation models are actively being used for remote sensing applications and several different algorithms have recently been proposed for various modalities and tasks. However, this study focuses on a foundation model used for SAR ATR (i.e., SAR image-based target classification and object detection). In the following sections, we introduce recent developments in remote sensing foundation models.
II-A Foundation models for remote sensing
@@
Remote sensing foundation models [44] have received widespread attention in recent years and have achieved effective learning across various modalities and tasks. Many of these studies have used existing large-scale pre-training datasets or have collected large quantities of samples from different sources. Model backbones have been established by improving attention mechanisms, positional encoding, and other aspects used to enhance the perception of complex spatial information. MIM has also been used to learn spatial-temporal contextual information, while contrast learning has been applied to multi-modal learning.
@@
SatMAE [45] involves a novel masking strategy with the temporal and spectral positional encoding used for multi-spectral and temporal images. SatMAE has also achieved excellent performance in scene classification and semantic segmentation tasks using a new dataset (fMoW Sentinel) that includes 13 different frequency bands. RVSA [46] improves a pre-training ViT backbone using a rotated variable-size window attention method for arbitrarily oriented objects. This work demonstrates the importance of learning complex spatial contextual relationships for targets in remote sensing images. RingMo [47] utilizes a patch incomplete mask strategy for dense or small objects and a self-constructed set of 2 million images, proving effective in a variety of tasks. It shows the potential of large-scale pre-training. RingMo-Sense [48] offers a three-branch network and a masking strategy for modeling spatio-temporal interactions in temporal images. CMID [49] combines contrastive learning and masked image modeling to learn global semantic information and local spatial information. This work also shows the importance of learning the diversity of contextual relationships in remote-sensing images. GFM [50] focuses on the differences between natural and remote sensing images and employs a multi-object continual pre-training approach to leverage information from both. It shows that detailed information in natural images can complement remote sensing images well. DiffusionSat [51] is the first remote sensing generative model to employ geographic information embedding in stable diffusion. Scale-MAE [52] reconstructs images at different frequencies with improved positional encoding for ViT. FG-MAE [43] employs various hand-designed features to replace original pixels in the MIM, thereby improving the feature quality. Similarly, SMLFR [53] uses a low-pass filter to eliminate high-frequency information from the image pixels. These studies show the necessity of SSL high-quality guide signals. As such, our work focuses on the issues of SAR image quality while considering the range and spatial resolution of remote sensing.
@@
A variety of multi-modal remote sensing FMs have also been developed, including SkySense [54] and OFA-Net [55]. SkySense [54] adopts a multi-granularity contrastive learning method to learn representations for different modalities. A GEO-context prototype has also been applied to embed geographical contextual information. OFA-Net [55] employs a shared transformer backbone for multiple modalities. In addition, vision-language models [56], such as EarthGPT [57], SkyEyeGPT [58], and LHRS-Bot [59] incorporate large language models into various remote sensing image modalities. However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only contains 10,554 SAR ship images, much less than the 84,838 infrared or 907,945 optical images. As a result, relatively little research is available on SAR FMs, due to the scarcity of high-resolution SAR target datasets (see Table I). This study explores visual FMs based on SAR images, using target recognition to promote research in this field.
II-B Related SSL in SAR
SSL for SAR ATR has been investigated in multiple studies. Early SSL was often used as a regularization loss for classification tasks, as discussed below. RotANet [60] predicts the rotational patterns of MSTAR vehicle targets by capturing azimuthal features for classification tasks. UACL [61] combines data augmentation and adversarial samples for contrastive learning, to improve model robustness to various adversarial attacks. PGIL [6] employs contrastive learning between complex SAR image sub-frequency features and deep amplitude image features, incorporating physical knowledge into classification tasks. SSL has also been used recently in model pre-training and fine-tuning frameworks. BIDFC [19] proposes weakly contrastive learning for pre-training in fine-grained vehicle datasets (MSTAR) and applied Gaussian noise data augmentation to simulate SAR image noise. TSCL [42] applies SAR image pre-processing prior to data augmentation in contrastive learning. FG-MAE [43] discusses different hand-crafted features for use with multi-spectral and SAR images and applies HOG features to SAR. In our previous studies, SAR-JEPA [39] SAR-JEPA [39] applies local reconstruction and multi-scale gradient features to collect target spatial signatures better. MSFA [21] proposes a multi-stage process with a filter augmentation pre-training framework for use in large-scale RGB and SAR data detection. However, these studies have only explored classification or detection tasks on a small number of datasets. Inspired by these previous works, this study aims to systematically investigate the construction of a SAR ATR foundation model for various ATR datasets via SSL.
Our Insights - These studies have demonstrated that SSL can achieve performance improvements across multiple categories [39] and tasks [39, 21], which may be comparable to the performance of specially designed supervised methods [19, 21]. However, it still lacks a foundation model for various SAR ATR applications. Besides, there is a lack of a pre-training and evaluation benchmark that contains images from classification and detection scenes. This inspired us to conduct systematic research into foundation models for general SAR target recognition, specifically with big data. We first extended the pre-dataset using different classification and detection tasks and scenarios (such as globally inland, marine, harbors, cities, and airports) based on existing research. A suitable model backbone is then discussed for the small target characteristics of remote sensing images. Since SSL requires high-quality guide signals from SAR images under the influence of noise, we applied two-step pre-training. Finally, we comprehensively evaluated the performance of the foundation models.
III Approach
We aim to construct a foundation model for general ATR from large-scale SAR images via an SSL method. As described above, increasing SAR datasets and SSL studies have inspired us to develop a foundation model for SAR ATR. This approach centers around pre-training datasets, model backbones, SSL methods, and evaluation tasks, to provide a systematic benchmark for SAR ATR foundation models.
@@ III Approach
TABLE I: A description of the SARatrX data, which included 14 open-source SAR datasets used for pre-training. Large SAR imagery in the detection datasets contained more targets and scene types than the annotations. Some large images were cropped to increase the number of available training samples. # Img.: Number of images. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Pol.: Polarization.
@@
Dataset Year # Imgs. Img. Size # Targets # Scenes Res. (m) Band Pol. Description
@@
III-A Establishing a Diverse Dataset for Pretraining
@@
Previous research primarily employed MSTAR [18] as a pretraining dataset. While MSTAR provides high-quality vehicle targets, it only contains a few thousand commonly used samples. The images also suffer from background bias caused by a single imaging scene [73]. In contrast, the ImageNet-1K pre-training set contains 1.4 million images with different categories and scenes. Since diverse target, scene, and sensor conditions constitute a large data sampling space in real-world scenarios, constructing a large pre-training dataset for the foundation model is central.
@@
The increasing availability of SAR target datasets as seen in Table. I, is a primary motivation for achieving this goal. Although SAR images are expensive and no single dataset contains all popular target categories or imaging conditions, collecting target samples from various open-source datasets can still provide a pre-training set with distinct categories, scenes, and sensors. As such, we constructed a new pre-training dataset, SARDet-180K, consisting of 186,600 SAR target samples from 14 open-source [^1] In this paper, we did not consider target datasets that are not open source or datasets for SAR terrain classification. Furthermore, we chose the typical open-source ship dataset as pre-training to avoid too many marine scenarios, and other open-source ship datasets about 28,749 images were not added, including FUSAR-Ship [74], LS-SSDD [75], DSSDD [76], HRSID [63], SRSDD [77], and RSDD [78]. SAR target datasets’ all images, as described in Table I. This set aims, to the extent possible, to include common target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of varying resolutions and bands).
III-B Model Architecture
@@
Two model backbones were considered for SAR target recognition. The first utilized a vision transformer (ViT) [79], commonly used in SSL, offering good scalability of model parameters. The second architecture employed ConvNeXt-V2 [80], which offered the same scalability as ViT but maintained the efficiency of a convolutional neural network. In addition to the scalability of model parameters, remote sensing tasks must also consider image properties. For example, SAR targets typically exhibit a small foreground and a dynamic context range. Our previous study, MSFA [21], demonstrated that Swin transformers could outperform ViTs with a hierarchical structure, yet are unsuitable when using drop patches in MIM to preserve computing resources. Therefore, we finally considered a variant of ViT, a hierarchical vision transformer (HiViT) [41], which improved the input spatial resolution and retained ViT properties for MIM.
Figure 2: The pre-training process included two steps. The first involved performing MIM on ImageNet data to obtain better initialization weights for model diversity, as shown in Fig. 4 (c). The open-source pre-training weights used in step 1 could easily be obtained, thereby reducing the pre-training time. The second involved performing MIM on SAR images with high-quality guide signals. Multi-scale gradient features were used to suppress speckle noise and extract target edges.
III-C Proposed Pretraining
MIM was used as a pretext task, and masked autoencoders (MAE) [81] were employed to drop patches and preserve computational resources. MIM can help foundation models achieve SAR image interpretation by recognizing contextual relationships in objects (a key point when applying MIM). SAR utilizes a type of coherent imaging, which involves speckle noise that can interfere with a pretext task. As such, SARatrX employs two pretraining steps to construct a foundation model, as shown in Fig. 2. The first step provides ImageNet weight to avoid the interference of SAR speckle noise in the early stages of the second pretraining step. Moreover, we use multiscale gradient features to suppress speckle noise throughout the second SAR pretraining step.
@@ III-C Proposed Pretraining
The first step involves performing MIM with ImageNet data to obtain better initialization weights because visible light images contain more and bettter signtures compared to SAR images. We simplified the multistage pretraining of MSFA, which performs SSL on ImageNet with a backbone, detection task pretraining on DOTA with a whole framework, and detection task finetuning on SAR images. SARatrX uses the pre-training weights from ImageNet as initialization weights for the SAR pre-training step. This approach enhances the diversity of attention during SAR pre-training, as shown in Fig. 4. In contrast, random initialization leads to the convergence of attention toward the same pattern as in SAR pre-training with MAE. The ImageNet pre-training backbone weights were obtained from an open source, thus reducing pre-training time significantly. This process of using pre-trained ImageNet weights is referred to as SSL-ImageNet & SAR.
@@
The second step involves performing MIM with SAR images. As mentioned previously, SAR image noise is a challenging problem that has been investigated in FG-MAR, SAR-JEPA, and MSFA, which have discussed features such as CannyEdge [82], HOG [83], Haar-like [84], SAR-HOG [85], and SAR-SIFT [86]. Different feature combinations can be used to achieve the best results [21], but the simplest gradient features follow our previous SAR-JEPA approach to avoid excessive runtimes during complex feature selection. As such, multi-scale gradient features (MGFs) [39] were used to suppress speckle noise and extract target shapes.
Multi-scale gradient feature - The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in SAR images. This implies that speckle can cause gradient calculations to produce false points in strong target regions. Previous studies [87, 88] have demonstrated that a computing ratio is suitable for multiplicative noise. In this study, MGF employed a gradient by ratio [85, 86] to obtain the relevant gradient features $G_{\text{m}}$:
$R_{i}$ $=\frac{M_{1}(i)}{M_{2}(i)},$ (1)
$G_{\text{H}}$ $=log(R_{1}),$ (2)
$G_{\text{V}}$ $=log(R_{3}),$ (3)
where $R_{i}$ denotes the average ratio in different directions. $M_{1}(i)$ and $M_{2}(i)$ denote area averages on opposite sides of a current pixel along the direction $i$, where $i=1$ represents the horizontal direction and $i=3$ indicates the vertical direction. These area averages can be calculated from an input image and four fixed convolution kernels (see Fig. 2). Eqs. 2 and 3 then use logarithms to perform the vertical gradient calculation [86], where $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient.
$G_{\text{m}}$ $=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}},$ (4)
MGF $=\text{concat}(G_{\text{m1}},G_{\text{m2}},G_{\text{m3}}),$ (5)
Due to the dynamic range required for various targets in remote sensing [89], MGF is constructed with convolutional kernels of different sizes. We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1 in order to calculate the average ratio in four different directions.
III-D Evaluation with Recognition Tasks
Fine-grained classification datasets comprised of vehicles, ships, and aircraft were merged to form a new SAR classification dataset called SAR-VSA (Vehicles, Ships, and Aircraft). This set was used to compare SSL model performance with the few-shot settings described in Sec. IV. We then report the SARatrX results for existing classification and detection settings, datasets, and algorithms in Sec. V.
IV SARatrX Experiments
We first performed SSL on the pre-training dataset without label information and then fine-tuned the pre-trained model on a classification dataset using few-shot classification tasks and linear probing settings to analyze improvements made to SARatrX. We also discuss the scalability of the proposed technique. Pre-training was first performed on eight NVIDIA RTX3090 GPUs. The SAR pre-training SARDet-180K dataset consisted of 14 SAR datasets (see Table I). Specifically, a few-shot SAR classification dataset, SAR-VSA, included 25 fine-grained targets from three SAR datasets (MSTAR [25], FUSAR-Ship [29], and SAR-ACD [29]). It is difficult to ensure training convergence by fine-tuning whole model parameters in small sample cases. As such, we used linear probing [81], which included a batch normalization layer, to adjust for differences in the statistical data properties and reduce the number of fine-tuning parameters. Detailed settings can be found in Appendix A.
@@ IV SARatrX Experiments
TABLE II: The results for a classification dataset containing 25 categories. Linear probing was performed during few-shot SAR target classification. We recommend using pre-trained ImageNet weights during initialization before pre-training the SAR datasets. We also recommend the HiViT model backbone for recognizing small targets in SAR images. Average accuracy was employed as the classification metric. Bolded text indicates the best result, while underlined text is the next best result. SL: supervised learning. SSL: self-supervised learning.
@@
Settings Dataset Method 5 10 20
@@
TABLE III: Comparisons of various target features for MIM. Many target features were found to be unsuitable for the multiplicative speckle noise in SAR images. This result inspired us to pursue different gradient features for SAR SSL. The pre-training settings in SSL-SAR and the model backbones formed the base version, while average velocity served as the classification metric. Bolded text indicates the best result, while underlined text is the next best result.
@@
IV-A Comparison of Model Backbones
Table II compares different model backbones used for SAR ATR, including ConvNeXt-V2 [80], ViT [79], and HiViT [41]. ConvNeXt-V2 and ViT represent the two primary architectures: a CNN and a transformer. Specifically, HiViT combines ViT with a Swin transformer [92]. The results indicated that ViT outperformed ConvNeXt-V2. On the one hand, ViT is more flexible than ConvNeXt-V2 for learning contextual information in SAR images. On the other hand, multiple downsampling steps in ConvNeXt-V2 resulted in the loss of small targets, while ViT maintained the same spatial resolution in different layers. HiViT, a visual transformer with ViT’s flexibility and hierarchical representation, outperformed ViT and ConvNeXt-V2 in our self-supervised experiments. In particular, HiViT employed small (4 × 4) input patches, capturing small target features well. Fig. 4 demonstrates that HiViT also offered a superior variable attention distance compared to ViT, due to the small target information common in remote sensing. Therefore, we use HiViT as backbone of our SARatrX.
IV-B Strategy of two-step pre-training
Here, we discuss the two-step pre-training strategy included to make full use of available model weights and SAR datasets. Table II includes four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR. SL-ImageNet was pre-trained on ImageNet using supervised learning [^2] We used open-source weights from GitHub to conduct the experiments. Supervised weights in ConvNeXt-V2 and HiViT were obtained with supervised fine-tuning after SSL.; SSL-ImageNet was pre-trained on ImageNet from scratch using SSL; SSL-SAR was pre-trained from scratch on our SAR pre-training dataset; SSL-ImageNet & SAR pre-trained the model on a SAR dataset based on initialized weights from SSL-ImageNet.
Notice the additional supervised information introduced by SL-ImageNet did not necessarily improve SAR ATR performance (e.g., the linear probing performance of SL-ImageNet for HiVit was lower than that of SSL-ImageNet). SSL-SAR achieved better results than SSL-ImageNet using less data (12%), reflecting large differences in target features between the two images. However, ImageNet pre-training weights did provide a good initialization for lower features, such as shape and texture in SSL, with visible spectral remote sensing [93] and medical images [94]. Our experiments also confirmed that using SSL-ImageNet as initialization weights improved the pre-training performance of SAR images (see Table II) and attention diversity (see Fig. 4). As such, SARatrX employed the SSL-ImageNet & SAR settings to complement the richness of pre-training.
IV-C Design of target signals for SAR images